THE FAR-INFRARED SPECTRUM OF ClNO_{2} (NITRYL CHLORIDE)

\underline{J. ORPHAL}, Institute of Environmental Physics, University of Bremen, P. O. Box 330440, 28334 Bremen, Germany; M. MORILLON-CHAPEY, LPMA,CNRS and University of Paris-Sud, Bat. 350, 91405 Orsay Cedex, France; S. KLEE, G. CH. MELLAU, and M. WINNEWISSER, Institute of Physical Chemistry, University of Giessen, H.-Buff-Ring 58, 35392 Giessen, Germany.

ClNO_{2} is a molecule of potential stratospheric and tropospheric importance, formed by heterogeneous reactions on polar stratospheric clouds and sea-salt particles. In 1994, high-resolution infrared spectra of ClNO_{2} were measured and analyzed in the 500–2500 cm^{−1} range^{a}. However, up to now, no high-resolution absorption spectra of the ν_{3} and ν_{5} fundamental bands of ClNO_{2} in the far-infrared region were reported, except for microwave observation of a few transitions in the excited vibrational states^{b}. In the present study, absorption spectra in 300–500 cm^{−1} range were recorded at resolutions of up to 0.0017 cm^{−1}, using the Bruker IFS-120HR FTS at University of Giessen. Using these spectra, nearly 2000 transitions of the a-type ν_{3} band around 370 cm^{−1} were assigned. The b-type ν_{5} band around 411 cm^{−1} is too weak to be observed, in agreement with ab-initio predictions. However, the (ν_{2} − ν_{3}) a-type difference band was observed and could be used to confirm rotational constants of both vibrational states.

aJ. Orphal, Ph.D. dissertation, University of Paris-Sud (1995)

bY. Morino and T. Tanaka, J. Mol. Spectrosc. 16, 179 (1965)